User Tools

Site Tools


lbaops:lbamar2007:v190htilog

07:36 UT Second LO freq. for the second channel has been corrected from 430MHz to 462 MHz.

lbaops/lbamar2007/v190htilog.txt · Last modified: 2015/12/18 16:38 (external edit)